Write the point-slope form of the line satisfying the given conditions. Then use the point-slope form of the equation to write the slope-intercept form of the equation.
Slope = 6, passing through (-7,4)

Answer by ankor@dixie-net.com(22740)

Write the point-slope form of the line satisfying the given conditions.
Then use the point-slope form of the equation to write the slope-intercept form of the equation.
Slope = 6, passing through (-7,4)
:
mx + b = y
where
x = -7
y = 4
m = 6
Find b
6(-7) + b = 4
-42 + b = 4
b = 4 + 42
b = 46
The equation
y = 6x + 46
:
:
Confirm this for yourself, replace x with -7, find y